OpenStudy (anonymous):

How would one get the x-intercepts of these exponentail graphs? y = 2 * 3^x + 1 and y = 3^-x -1

OpenStudy (anonymous):

Isolate 3^x then use log, base-3.

OpenStudy (anonymous):

uh... without a calculator?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Do you want exact values or approximations?

OpenStudy (anonymous):

the exact values.

OpenStudy (anonymous):

The first one will give you problems because of taking log of a negative number, so the solution there will be complex (imaginary).

OpenStudy (anonymous):

The second one is easy because the log of 1 = 0.

OpenStudy (campbell_st):

well look at the 1st graph... an exponential graph y = a^x has a horizontal asymptote at x = 0 so y = a^x + 1 has the asymptote at x = 1 so the graph will not have an x - intercept.

OpenStudy (anonymous):

That makes sense, Campbell_st. We havent been taught logarithms, Cliffsedge.

OpenStudy (campbell_st):

the 2nd graph y = 3^(-x) - 1 has a horizontal asymptote at y = -1 y = 3^{-x} would cut the y axis at y = 1 since the graph is moved down 1 the y intercept is y = 0 which means the x intercept is x = 0 with a horizontal asymptote at y = -1